What force should be applied so that a 10kg block accelerates at 5.0m/s2 up a 20° inclined plane as it is pulled from the top? The object's coefficient of kinetic friction is 0.10.


1
Expert's answer
2021-12-20T10:28:09-0500

The force is


"F=m(a+\mu g\sin\theta)=53.4\text{ N}."